Estimating the Cost of Debt Capital

Kellogg Company manufactures cereal and other convenience food under its many well-known brands such as Kellogg’s®, Keebler®, and Cheez-It®. The company, with over $13.5 billion in annual sales worldwide, partially finances its operation through the issuance of debt. At the beginning of its 2015 fiscal year, it had $6.7 billion in total debt. At the end of fiscal year 2015, its total debt had increased to $6.8 billion. Its fiscal 2015 interest expense was $418 million, and its assumed statutory tax rate was 37%.

a. Compute the company’s average pretax borrowing cost. (Hint: Use the average amount of debt as the denominator in the computation.)

Round your answer to one decimal place (ex: 0.0345 = 3.5%).

b. Assume that the book value of its debt equals its market value. Then, estimate the company’s cost of debt capital.

Round your answer to one decimal place (ex: 0.0345 = 3.5%).
Answer #1

Debt is used to represent any security which bears a fixed financial charge. The cost of debt refers to the cost associated with using fixed interest bearing securities

To calculate pre-tax cost of debt, take the sum total of debt-related interest payments divided by the total amount of debt taken on for the year.

Annual sales - $13.5 billion

Total outstanding Debt - $6.8 billion

Interest expenses- $418 million

A) Average Pre tax borrowing cost -To calculate pre-tax cost of debt, take the sum total of debt-related interest payments divided by the total amount of debt taken on for the year.

=Total interest payments/ Average outstanding debt

= $418 million/ $6.75 billion

= 0.0619 or 6.19 %

Average outstanding debt = ($6.7+ $6.8)/2 =$ 6.75 billion

B) Cost of debt capital - it must be noted that the interest payable on debt capital is tax deductible expense. Therefore company calculates cost of debt after tax which is termed as ‘ post tax cost of debt’

After tax cost of debt = Pretax cost of debt * (1- tax rate)

= 6.19 *(1-0.37) = 3.89%